SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 1 Dec 2003

Status Report From: NASA HQ
Posted: Monday, December 1, 2003

image All ISS systems continue to function nominally, except as noted previously or below.  Week 6 is underway for Expedition 8.

Flight Engineer Kaleri made preparations for his first experiment session with the Russian/German Plasma Crystal-3 (PK-3) payload, planned for tomorrow.  Preps consisted of activating the evacuation turbopump, checking for leaks and tagging up with ground specialists, then starting the evacuation of the vent lines and vacuum chamber (ZB).  Later, the experiment was terminated and the turbopump was to be deactivated shortly before sleep time.  [The experiment will be performed on plasma, i.e., fine particles, charged and "excited" by RF/radio frequency power, inside the evacuated work chamber.]

CDR/SO Michael Foale worked with the CSA-CP (compound specific analyzer-combustion products). [He first conducted a "zero" calibration (i.e., re-zeroing gas sensors), then marked the unusable O2 sensor on the backup CSA-CP as "failed", and completed a data take with the CSA-CP, calling down O2, CO, HCl, HCN and battery status values for both units, which he then deactivated.]

An agreement between the two MCCs (Houston & Moscow) will obviate the need for timelining daily CSA-CP O2 readings for Mike Foale and ensure that everybody has the data that they need.

The monthly PEP (personal emergency provisions) inspection procedure has been updated for easier use by the crew, starting with the upcoming inspection on 12/5. [The procedure involves verification that PFEs (portable fire extinguishers), PBAs (portable breathing assemblies), QDMAs (quick-don mask assemblies) and EHTKs (extension hose/tee kits) are free of damage, to ensure their functionality.  It will not be necessary to remove any PEPs from their locker unless obvious damage is discovered during the inspection.]

Mike Foale was asked to locate a missing EHTK for use with the PBA.  [During the last few emergency equipment inspections, the EHTK assigned to the Node starboard emergency equipment locker could not be located.  Since it is required in the event of an emergency, Mike was to search for and co-locate it with the Node PBAs, along with spare pre-breathe hose assembly (PHA) equipment.]

Mike also cleaned the QDMAs, which were noticed by the ground during a recent routine Ku-band video downlink to be dirty.

After the failure of the transfer of Foale's personal montage file on an HRF (human research facility) flash card to the HRF PC into a specific directory for the upcoming experiment session of FOOT (Foot/Ground Reaction Forces during Space Flight) on 11/28 (PC not recognizing the ADAS flash card), instructions were uplinked for Mike to troubleshoot and correct the problem today.

Kaleri completed a 2-hr. task in the FGB (Funktsionalnyi-Grusovoi Blok), removing and replacing the no. 2 current converter control device (BUPT-2) of battery #6 with a spare unit, discarding the removed component.

The CDR initiated another maintenance charging/discharging cycle on EMU/spacesuit batteries in the Airlock's BSA (battery stowage assembly), today on batteries #2029 and #2030.  [The charging will take about 24 hrs and will be followed by discharge.  Helmet light and PGT (pistol grip tool) batteries were not charged at this time and had to be removed from the BSA beforehand.]

Mike also replaced the batteries of two IWIS RSUs (internal wireless instrumentation system/remote sensor units), #1026 and #1027.

The crew completed another EPO (Educational Payload Operations) activity.  Before starting the demonstrations, Mike and Sasha set up the camcorder for Sasha to record and downlink the sessions via Ku- and S-band in real-time for taping on the ground, to be used in NASA educational products, on websites, in schools, on TV, etc. [Today's two EPO activities explained (1) the ISS flight attitudes and how the station is controlled and oriented with gyroscopes and thrusters, and (2) a simple demonstration of Bernoulli's Principle.]

The CDR conducted the daily routine maintenance of SOZh life support systems (including toilet facility, food containers, water containers and solid waste containers) and the regular routine checkup of Increment 8 payloads running in the Lab, while the FE attended to the preparation of the daily IMS (inventory management system) update file.   The crew worked out according to their regular daily physical exercise program of 2.5 hrs on the TVIS treadmill (the modified use of which has been approved for two weeks of further data collection), RED exerciser, CEVIS cycle ergometer and VELO bike (with load trainer).

Foale, assisted by Kaleri, performed the periodic on-orbit calibration of the two RED Flexpack canisters #1001 and #1003.

The Russian VELO ergometer continues to be operating (after responding favorably to a "kick" last week, 11/24), but TsUP/Moscow has scheduled a troubleshooting/IFM (inflight maintenance) task tomorrow for possible identification of its prior malfunction.

At 12:17pm EST, MCC-H switched S-band transmissions for 15 min. from the regular S/G (space-to-ground) rate to low data rate (LDR).  [This routine test is done four times per year to make sure LDR is available if needed in an emergency.]

Plans are being discussed to introduce the NGL (next generation laptop) also in the Russian segment (RS).  [The crew has repeatedly commented favorably on the performance of the NGLs in the USOS (U.S. segment).  The NGL is an IBM ThinkPad A31p running on a Pentium IV/2 GHz microprocessor.]

Last week (11/28), Russia's State Duma approved the fourth and final reading of the draft federal budget for 2004, which would increase funding for all space programs (including ISS) by 6.3 billion rubles over FY 2003 spending (i.e., by ~$207 million to ~$470 million). [Approval by the Federation Council (Upper House of Parliament) is now needed before it can be signed by Pres. Putin.]   U.S. and Russian Segment Status (as of 1:30pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 24A.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) -- 742; temperature (deg C) -- 25.7; ppO2 (mmHg) -- 158.3; ppCO2 (mmHg) -- 4.6.
  • SM Transfer Compartment:  Pressure (mmHg) -- 761; temperature (deg C) -- 19.6.
  • FGB Cabin:  Pressure (mmHg) -- 756; temperature (deg C) -- 22.3.
  • Node:  Pressure (mmHg) -- 754.07; temperature (deg C) -- 22.3 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • U.S. Lab:  Pressure (mmHg) -- 755.93; temperature (deg C) -- 25.0; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) -- 756.03; temperature (deg C) -- 24.0; shell heater temp (deg C) -- 23.2, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • PMA-1:  Shell heater temp (deg C) -- 22.7
  • PMA-2:  Shell heater temp (deg C) -- 20.4

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed (dual-angle) Position (non-suntracking). 
  • SM batteries:  Battery #8, formerly known as #7, is still in slot #8 for troubleshooting; all other batteries (7) are in "Partial Charge" mode.  
  • FGB batteries:  Battery #5 is off; all other batteries (5) are in "Partial Charge" mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational (all lanes reintegrated 11/5).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; string #3 dropped out 10/22.

Propulsion System:

  • Total propellant load available:3704 kg (8166 lb) as of 11/13 [SM(755) + FGB(2597) + Progress M(352) + Progress M-1(0)].  (Capability: SM -- 860 kg; FGB -- 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source -- U.S. SIGI-1 (GPS)
  • Attitude source -- U.S. SIGI-1 (GPS)
  • Angular rate source -- RGA-1

Flight Attitude:

  • LVLH  XVV (local vertical/local horizontal = "earth-fixed": z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:46am EST [= epoch]):

  • Mean altitude -- 373.0 km
  • Apogee -- 377.1 km
  • Perigee -- 369.0 km
  • Period -- 92.0 min.
  • Inclination (to Equator) --  51.63 deg
  • Eccentricity -- 0.0005966
  • Orbits per 24-hr. day -- 15.65
  • Mean altitude loss in last 24 hours -- 200 m
  • Revolutions since FGB/Zarya launch (Nov. '98)  -- 28715
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see
  •  http://www.hq.nasa.gov/osf/station/viewing/issvis.html

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter